10. Use the Midpoint Rule with n = 4 to approximate the area under the curve the interval (1,5). f(x) = V2 +6 on
Answer #1

h = 3.061 +4.650 +6.991 +9.855 f60) = 3+1 5] for n=4 b = b-a =y hel n ។ midpoint rule Area = Š 4(**-**33) = 1 f( xox*) sto H